OpenStudy (anonymous):

I need help understanding how to do this. I have a test tomorrow. I know I should know this but I am drawing blanks right now. 2/x+1-1/x=2x^2+x

OpenStudy (anonymous):

\[\frac{2}{x+1}-\frac{1}{x}=2x^2+x\] like that ?

OpenStudy (anonymous):

i hope not because i don't think this actually has any solutions at least not real solutions

OpenStudy (anonymous):

yes, ut the last one after the equal sign is 2/ x^2+x

OpenStudy (anonymous):

oooh \[\large \frac{2}{x+1}-\frac{1}{x}=\frac{2}{x^2+x}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

subtract on the left, the denominator will be \(x(x+1)=x^2+x\) same as on the right

OpenStudy (anonymous):

so far that is the same answer i have so far aswell x(x+1)

OpenStudy (anonymous):

\[\frac{2}{x+1}-\frac{1}{x}=\frac{2x-(x+1)}{x(x+1)}\] \[=\frac{x-1}{x(x+1)}\] now solve \[\frac{x-1}{x(x+1)}=\frac{2}{x(x+1)}\] by \[x-1=2\] etc

OpenStudy (anonymous):

ok thanks